# Gravitational collapse of a degenerate wormhole

### v1

@@ Abstract
The dynamics of a degenerate spherically symmetric wormhole in a vacuum is considered. An extension of the equivalence principle to matter-free objects that are the source of a gravitational field is proposed. Using the Klinkhamer metric as an example, it is shown that a degenerate wormhole is precisely such an object. Application of the extended equivalence principle reduces the radial dynamics of the Klinkhamer wormhole to the dynamics of the radial fall of a test particle in a Schwarzschild gravitational field. It is proven that any bound state of the traversable Klinkhamer wormhole eventually collapses into a non-traversable Einstein-Rosen wormhole. An estimate is presented showing that the traversable Klinkhamer wormhole, although non-stationary, is a long-lived state.
@@ 1 Introduction
Since the degenerate Einstein-Rosen metric describes an empty and static curved space, the corresponding wormhole is an example of a matter-free stationary source of a gravitational field.
@@
Recent works by Klinkhamer (Klinkhamer, 2022, 2023, 2025; Wang, 2023) suggest a return to degenerate wormholes, where it all began. The Klinkhamer wormhole metric is a direct generalization of the Einstein-Rosen metric. The generalization is that the radius $b$ of a Klinkhamer wormhole can be larger than the gravitational radius, which ensures traversability. Constructed in complete analogy with the Einstein-Rosen metric, the Klinkhamer metric is also degenerate. Therefore, it should also be described not by the standard Einstein equations but by regularized equations—for the same reason that Einstein and Rosen did so in 1935. Like the Einstein-Rosen wormhole, the Klinkhamer wormhole is a matter-free source of a gravitational field, i.e. it possesses gravitational mass. By the equivalence principle, this must also imply that the Klinkhamer wormhole possesses an inertial mass.
@@
Thus, the subject of this work is the dynamics of a degenerate wormhole using the Klinkhamer wormhole as an example.
@@
The paper is organized as follows. In the second section, we consider the regularized Einstein equations and the Klinkhamer metric as one possible stationary solution of this system of equations. In the third section, we generalize the stationary Klinkhamer metric to a general non-stationary solution of the regularized equations, describing arbitrary wormhole dynamics within the framework of spherical symmetry. In the fourth section, we extend the equivalence principle to matter-free objects, an example of which is a degenerate wormhole. In the next, fifth section, we directly apply the extended equivalence principle, which allows us to reduce the dynamics of a spherically symmetric wormhole to the dynamics of a single test particle in a Schwarzschild gravitational field. In the sixth section, we present a well-known solution to the problem of the radial fall of a test particle in a Schwarzschild gravitational field. This allows us to obtain a unique solution to the problem of the dynamics of a degenerate spherical wormhole under given initial conditions in the seventh section, in particular, a complete solution to the gravitational collapse problem. The collapse results in the transformation of the Klinkhamer wormhole into an Einstein-Rosen wormhole. In the final, eighth part, we summarize and draw conclusions.
@@ 2 Stationary solution
We begin by presenting the Klinkhamer metric describing a degenerate spherical wormhole in a vacuum. We choose the Schwarzschild metric as the initial one [^1] The relativistic system of units is used, in which the speed of light $c$ and the gravitational constant $G$ are equal to 1, as well as the usual spherical coordinates $(r,\theta,\varphi)$, $d\Omega^{2}=sin^{2}\theta d\varphi^{2}+d\theta^{2}$ is an element of solid angle, $t$ is time, $M$ is the mass of the source of the gravitational field.
@@
which transforms the original one-sheeted space into the two-sheeted space of the wormhole. The new variable $l$ covers all values from -$\infty$ to $\infty$, with positive values $l>0$ corresponding to the first sheet, negative values $l<0$ to the second sheet of the two-sheeted space, and the value $l=0$ to the transition surface between them (i.e., the spherical throat $r=b$). The constant parameter $b\geq 2M$ represents the radius of the throat. The topological dressing transforms the original solution (1) of the standard Einstein equations (2) into a degenerate Klinkhamer metric (Klinkhamer, 2022)
@@ 3 Non-stationary solution
4 Extended equivalence principle
When describing the dynamics of a degenerate wormhole under the influence of its own gravitational field, it is important to understand the source of this field. Consider this using the example of the Klinkhamer wormhole.
@@ 4 Extended equivalence principle
To do this, let us compare the Schwarzschild metric (1) and the Klinkhamer metric (7), written in the same $r$-coordinates. Both metrics contain the same mass parameter $M$, but its meaning in these two metrics is different. In the Schwarzschild metric, it is the mass of the centrally symmetric distribution of matter in 3D space. The Klinkhamer metric is a solution of the regularized Einstein equations (5) in the entire two-sheeted space, so matter is absent here. In this case, the source of the field is not matter, but the 2D surface of the two-sheeted space—its spherical throat $r=b$. Due to the symmetry of the two-sheet metric (4) with respect to the permutation of the first and second sheets (i.e., the change in the sign of the two-sheet coordinate $l$), this gravitational field is the same on both sheets.
Thus, the wormhole turns out to be an object endowed with a gravitational mass $M$. Does this object possess inertia? Without an answer to this question, it is impossible to construct the dynamics of a wormhole. To answer this question, we must turn to the equivalence principle, which postulates ”the equality (in suitable units) of the inertial and gravitational masses, regardless of the nature of the body(Misner et al., 1973). Usually, the term ”body” implies that matter fills a region of 3D space. In the case of a wormhole, we are dealing with a different object—a compact gravitating 2D surface free of matter. Therefore, in this case, the equivalence principle must be expanded to ”the equality (in suitable units) of the inertial and gravitational masses, regardless of the nature and dimensionality of the object.” This allows us to include not only matter but also objects that are self-consistent solutions of Einstein’s equations in a vacuum. We will call this formulation of the equivalence principle ”extended”.
@@
How acceptable is such an extension? A necessary condition for its acceptability is the absence of internal contradictions and the ”integration” of its consequences into an already accepted and tested system of concepts and physically measurable quantities.
In considering the radial dynamics of the Klinkhamer wormhole, we will proceed from this extended interpretation of the equivalence principle. In any case, this cannot lead to contradictions with traditional general relativity, which is based on the standard equivalence principle. In studying the radial dynamics of the wormhole, we will demonstrate that this interpretation leads to natural results that are consistent with our understanding and are entirely analogous to the conventional picture of gravitational collapse.
@@ 5 Gravitational collapse and the extended equivalence principle
In this section, we show that a direct application of the extended equivalence principle reduces the field problem of the radial dynamics of a centrally symmetric wormhole (here, the Klinkhamer wormhole) to a single-particle problem of the motion of a test particle in a given gravitational field.
@@
By virtue of the extended equivalence principle, the position of the test particle relative to the wormhole surface should remain unchanged during the collapse process (the ”falling elevator effect”). This means that $b(t)$ must be identical to the law of radial motion of a particle $r(t)$ in a given gravitational field of the wormhole (see Fig. 1).
Thus, the problem of wormhole collapse is reduced to the problem of free fall of a test particle located near the surface of a collapsing wormhole.
Figure 1: By virtue of the extended equivalence principle, the collapse of a degenerate wormhole of mass $M$ is synchronous with the free fall of a test particle of mass $m$ located near the surface of the wormhole.
@@ 6 Free fall of a test particle
7 Radial dynamics of Klinkhamer’s wormhole
@@
We have shown that, based on the extended equivalence principle, the dependence of the wormhole radius on time $b(t)$ should be identical to the dependence $r(t)$ for a test particle that falls freely. This means that the equation for the radius of the wormhole $b(t)$ is obtained from equation (16) by replacing $r(t)$ with $b(t)$:
$\frac{db}{dt}=\pm\left(1-\frac{2M}{b}\right)\left[1-\left(\frac{M}{E}\right)^{2}\left(1-\frac{2M}{b}\right)\right]^{1/2}.$ (17)
@@
Here, the minus sign corresponds to the case of gravitational collapse (a decrease in the radius $b$), and the plus sign corresponds to the opposite case of inertial expansion (an increase in the radius $b$). Simultaneously, we replace the mass $m$ of the test particle with the mass $M$ of the wormhole and the total energy $\varepsilon$ of the test particle in a given gravitational field with the self-energy $E$ of the wormhole. The correctness of this substitution is verified by the Newtonian limit, when the collapse velocity is small $db/dt=v\ll 1$, and the wormhole radius $b$ is much larger than the gravitational radius $2M/b\ll 1$. Then, equation (17) leads to a non-relativistic expression for the total energy in a central gravitational field:
$E=M+\frac{Mv^{2}}{2}-\frac{M^{2}}{b}=\text{const},$ (18)
as expected [^4] A small but revealing subtlety: Newton’s expression (18) contains the potential energy of a spherical wormhole as $M^{2}/b$, while the Newtonian potential energy of a hollow sphere of radius $b$ is half as large. The reason for this is that for a hollow sphere in an ordinary one-sheeted space, this energy can be obtained by integrating the Newtonian energy density $\mathfrak{g}^{2}/(8\pi),\>\mathfrak{g}=M/r^{2}$ over the outer region $r>b$. In a two-sheeted space, the result of the same integration is twice as large, since it covers not one, but two sheets..
@@
In each of these three regimes, the Klinkhamer wormhole, although non-stationary, nevertheless remains traversable for a certain time. This time depends on the initial conditions and can be quite significant. For example, the collapse from a resting state with an initial radius $b(0)\equiv b_{0}\gg 2M$ takes a time $\tau$ equal to the fall time of a test particle from a distance $b_{0}$ in the gravitational field of a classical point source of mass $M$. This time is (Landau and Lifshitz, 1982)

### v2

@@ Abstract
The dynamics of a degenerate spherically symmetric wormhole in a vacuum is considered. An extension of the equivalence principle to matter-free objects that are the source of a gravitational field is proposed. Using the Klinkhamer metric as an example, it is shown that a degenerate wormhole is precisely such an object. Applying the equivalence principle to matter-free field sources reduces the radial dynamics of the Klinkhamer wormhole to the dynamics of the radial fall of a test particle in a Schwarzschild gravitational field. It is proven that any bound state of the traversable Klinkhamer wormhole eventually collapses into a non-traversable Einstein-Rosen wormhole. An estimate is presented showing that the traversable Klinkhamer wormhole, although non-stationary, is a long-lived state.
@@ 1 Introduction
Since the degenerate Einstein-Rosen metric describes an empty and static curved space, the corresponding wormhole is an example of a matter-free stationary source of a gravitational field. In their original paper on the Einstein-Rosen bridge, its authors pointed out that the degenerate, geodesically complete solution they constructed of the regularized Einstein equations itself ”describes the field-producing mass without requiring the introduction of any new field quantities” (Einstein and Rosen, 1935), Section 1.
@@
Recent works by Klinkhamer (Klinkhamer, 2023a, b, 2025; Wang, 2023) suggest a return to degenerate wormholes, where it all began. The Klinkhamer wormhole metric is a direct generalization of the Einstein-Rosen metric. The generalization is that the radius $b$ of a Klinkhamer wormhole can be larger than the gravitational radius, which ensures traversability. Constructed in complete analogy with the Einstein-Rosen metric, the Klinkhamer metric is also degenerate. Therefore, it should also be described not by the standard Einstein equations but by regularized equations—for the same reason that Einstein and Rosen did so in 1935.Analogously to the Einstein–Rosen wormhole, the Klinkhamer wormhole constitutes a vacuum (purely geometric) solution that possesses an effective gravitational mass $M$, without the need to invoke any additional matter sources.
@@
This paper examines the dynamics of a degenerate wormhole using the example of the degenerate Klinkhamer wormhole. Specifically, the process of its gravitational collapse will be considered.
Investigating the dynamical properties of the Klinkhamer wormhole necessitates an appropriate generalization of its original stationary metric to a broader class of nonstationary metrics, in such a way that the stationary configuration arises as a particular case. Describing the dynamics of a degenerate wormhole can be pursued through three possible approaches :
1) by employing exclusively the regularized Einstein field equations themselves;
2) by formulating and analyzing a reduced action, from which one derives the corresponding equations of motion for a suitably reduced set of dynamical variables;
3) by constructing a reduced description grounded not on an action principle but rather on the equivalence principle, which likewise leads to a system of equations governing a reduced set of variables.
Let us evaluate these three possibilities sequentially.
1) Restricting the analysis solely to the regularized Einstein field equations is insufficient and reveals a fundamental limitation of this approach. The underlying reason is that the local metric $g_{\mu\nu}$ by itself does not furnish a complete characterization of the wormhole. A full description is provided instead by the pair $(\mathcal{M},g_{\mu\nu})$, which, in addition to the local metric $g_{\mu\nu}$, incorporates the global topological properties of the manifold $\mathcal{M}$. As demonstrated in (Dimaschko, 2024), once this fact is taken into account, the application of the principle of least action yields, in a unified manner, both the Einstein field equations for $g_{\mu\nu}$ and separate evolution equations for the topological characteristics of the manifold $\mathcal{M}$. It is precisely this circumstance that gives rise to the aforementioned Feng paradox.
Specifically, in the context of the radial collapse of a degenerate wormhole, the sole topologically relevant degree of freedom is the time-dependent throat radius $b(t)$. The Einstein field equations are insensitive to the detailed temporal profile of $b(t)$; they impose only that the spacetime metric reduces to the Schwarzschild form in the exterior domain $r>b(t)$. As a result, the regularized Einstein equations, considered in isolation, are insufficient to capture the complete dynamical evolution of a degenerate wormhole.
2) Employing a reduced action would be a natural strategy in this context; however, it results in a highly intricate mathematical problem. The underlying reason is that the Einstein–Hilbert action functional, based on the scalar curvature, becomes ill-defined when the metric is degenerate. Consequently, it is necessary to resort to a polynomial action (Peres, 1963; Katanaev, 2006), which possesses a substantially more involved structure.
The formalism based on reduced variables is also well established within the thin-shell model, which arises as the zero-thickness limit of the Thorne framework as the layer thickness approaches zero Visser (1996). In this formulation, however, the spacetime metric remains nondegenerate and matter fields are present, thereby placing the Klinkhamer wormhole solution—characterized by a degenerate metric in the absence of matter—clearly outside scope of this model.
3) Owing to its spherical symmetry, the degenerate Klinkhamer wormhole possesses a single dynamical degree of freedom, namely the wormhole radius $b$. Under these circumstances, the field dynamics effectively reduce to a one-dimensional problem, rendering it natural to invoke the equivalence principle. This principle is conventionally formulated in the context of point-particle dynamics and, to date, has not been systematically employed to construct the dynamics of field configurations. The Klinkhamer wormhole thus provides the first known example of a spacetime metric that permits the direct application of the equivalence principle in this manner, thereby mapping a field-theoretic problem onto an equivalent single-particle problem.
This third approach is the one adopted in the present work. Its objective is to formulate the dynamical properties of a degenerate Klinkhamer wormhole on the basis of the equivalence principle. The proposed framework is predicated on extending the equivalence principle to field-theoretic (geometric) configurations that are free of matter sources. In the particular case of a degenerate wormhole, the central geometric object of interest is the wormhole throat.
@@
The paper is organized as follows. In the second section, we consider the regularized Einstein equations and the Klinkhamer metric as one possible stationary solution of this system of equations. In the third section, we generalize the stationary Klinkhamer metric to a general non-stationary solution of the regularized equations, describing arbitrary wormhole dynamics within the framework of spherical symmetry. In the fourth section, we extend the equivalence principle to matter-free objects, an example of which is a degenerate wormhole. In the next, fifth section, we directly apply the equivalence principle, which allows us to reduce the dynamics of a spherically symmetric wormhole to the dynamics of a single test particle in a Schwarzschild gravitational field. In the sixth section, we present a well-known solution to the problem of the radial fall of a test particle in a Schwarzschild gravitational field. In the seventh section, this allows us to obtain a unique solution to the problem of the dynamics of a degenerate spherical wormhole under given initial conditions, in particular, a complete solution to the gravitational collapse problem. The collapse results in the transformation of the Klinkhamer wormhole into an Einstein-Rosen wormhole. In the same section, we also examine the consistency between the equations of motion derived from the equivalence principle and the dynamical equations obtained in the Newtonian limit. In the final, eighth part, we summarize and draw conclusions.
@@ 2 Stationary solution
We begin by presenting the Klinkhamer metric describing a degenerate spherical wormhole in a vacuum. We choose the Schwarzschild metric as the initial one [^1] The relativistic system of units is used, in which the speed of light $c$ and the gravitational constant $G$ are equal to 1, as well as the usual spherical coordinates $(r,\theta,\varphi)$, $d\Omega^{2}=d\theta^{2}+sin^{2}\theta d\varphi^{2}$ is an element of solid angle, $t$ is time, $M$ is the mass of the source of the gravitational field.
@@
which transforms the original one-sheeted space into the two-sheeted space of the wormhole. The new variable $l$ covers all values from -$\infty$ to $\infty$, with positive values $l>0$ corresponding to the first sheet, negative values $l<0$ to the second sheet of the two-sheeted space, and the value $l=0$ to the transition surface between them (i.e., the spherical throat $r=b$). The constant parameter $b\geq 2M$ represents the radius of the throat. The topological dressing transforms the original solution (1) of the standard Einstein equations (2) into a degenerate Klinkhamer metric (Klinkhamer, 2023a)
@@ 3 Non-stationary solution
4 Extended interpretation of the equivalence principle
In the analysis of the dynamics of a degenerate wormhole subjected to its self-generated gravitational field, it is essential to identify and characterize the source of this field. This issue can be elucidated concretely by considering the specific example of the Klinkhamer wormhole.
@@ 4 Extended interpretation of the equivalence principle
To proceed, we compare the Schwarzschild metric (1) with the Klinkhamer metric (7), expressed in the same radial coordinate $r$. Both metrics involve the same mass parameter $M$, but its physical interpretation differs in the two cases. In the Schwarzschild metric, $M$ represents the mass of a spherically symmetric distribution of matter in three-dimensional space. By contrast, the Klinkhamer metric is an exact solution of the regularized Einstein equations (5) on the entire two-sheeted manifold, where no matter is present. In this latter case, the source of the gravitational field is not matter in the usual sense, but rather the two-dimensional surface associated with the two-sheeted geometry—specifically, its spherical throat at $r=b$. Owing to the symmetry of the two-sheet metric (4) under the exchange of the two sheets (equivalently, the change of sign of the two-sheet coordinate $l$), the resulting gravitational field is identical on both sheets.
Thus, the wormhole can be regarded as an object endowed with a gravitational mass $M$. A natural question then arises: does this object also possess inertia? Without addressing this issue, it is not possible to formulate a consistent dynamical description of a wormhole. To resolve this, we invoke the equivalence principle, which postulates
“the equality (in suitable units) of the inertial and gravitational masses, regardless of the nature of the body(Misner et al., 1973).
In conventional usage, the term “body” refers to a material distribution occupying a three-dimensional spatial region. In the wormhole case, however, the situation is fundamentally different: we are dealing with a compact, gravitating two-dimensional surface that is devoid of matter. Accordingly, the standard formulation of the equivalence principle must be generalized to assert
the equality (in suitable units) of the inertial and gravitational masses, regardless of the nature and dimensionality of the object.
@@
What criteria determine the admissibility of such an extension? A necessary condition for its acceptability is the absence of internal inconsistencies, as well as the successful incorporation of its implications into an already established and empirically tested system of theoretical concepts and operationally defined, physically measurable quantities.
In Section 7 of this paper, we show, that in the Newtonian limit, the extended formulation of the equivalence principle precisely corresponds to the principles of ordinary classical mechanics extended to a two-sheeted space. In this sense, it passes the test of the correspondence principle and allows for a natural extension to general relativity. Therefore, in our further consideration of the radial dynamics of the Klinkhamer wormhole, we will always proceed from this extended interpretation of the equivalence principle.
Thus, we regard the extended equivalence principle as an axiom, on the same footing as the standard equivalence principle in general relativity, and will next test it by consistency and correspondence arguments.
@@ 5 Gravitational collapse and the extended equivalence principle
In this section, we show that a direct application of the equivalence principle reduces the field problem of the radial dynamics of a spherically symmetric wormhole (here, the Klinkhamer wormhole) to a single-particle problem of the motion of a test particle in a given gravitational field.
@@
By virtue of the equivalence principle, the position of the test particle relative to the wormhole surface should remain unchanged during the collapse process (the ”falling elevator effect”). This means that $b(t)$ must be identical to the law of radial motion of a particle $r(t)$ in a given gravitational field of the wormhole (see Fig. 1).
Thus, the problem of wormhole collapse is reduced to the problem of free fall of a test particle located near the surface of a collapsing wormhole. [^4] In this case, the possibility of applying the equivalence principle is due to the absence of non-gravitational forces. In (Dimaschko, 2024), a charged wormhole was considered against the background of an electrovacuum. In this case, the equivalence principle was inapplicable due to the presence of non-gravitational forces. Therefore, to determine the equilibrium radius of the wormhole, a direct application of the principle of least action was required.
Figure 1: By virtue of the equivalence principle, the collapse of a degenerate wormhole of mass $M$ is synchronous with the free fall of a test particle of mass $m$ located near the surface of the wormhole.
@@ 6 Free fall of a test particle
7 Radial dynamics and self-energy of the wormhole
@@
We have shown that, based on the equivalence principle, the dependence of the wormhole radius on time $b(t)$ should be identical to the dependence $r(t)$ for a test particle that falls freely. This means that the equation for the radius of the wormhole $b(t)$ is obtained from equation (16) by replacing $r(t)$ with $b(t)$:
$\frac{db}{dt}=\pm\left(1-\frac{2M}{b}\right)\left[1-\left(\frac{M}{\mathcal{E}}\right)^{2}\left(1-\frac{2M}{b}\right)\right]^{1/2}.$ (17)
@@
Here, the minus sign corresponds to the case of gravitational collapse (a decrease in the radius $b$), and the plus sign corresponds to the opposite case of inertial expansion (an increase in the radius $b$). Simultaneously, we replace the mass $m$ of the test particle with the mass $M$ of the wormhole and the total energy $\varepsilon$ of the test particle in a given gravitational field with the ”self-energy” $\mathcal{E}$ of the wormhole.
In fact, the quantity $\mathcal{E}/M$, which appears in the new differential equation (17), inherits from the ratio $\varepsilon/m$, which enters the original differential equation (16), only its kinematic meaning—the meaning of the first integral of motion. This meaning, in complete analogy with (16), consists of nothing more than the conservation of a certain combination of the radius $b$ of the wormhole and the rate $db/dt$ of change of this radius:
$\left(1-\frac{2M}{b}\right)^{-1}\left[1-\left(1-\frac{2M}{b}\right)^{-1}\left(\frac{db}{dt}\right)^{2}\right]=\text{const}=\left(\frac{M}{\mathcal{E}}\right)^{2}.$ (18)
To clarify the independent dynamic meaning of the ”self-energy” $\mathcal{E}$, we will proceed in the same way as is customary to clarify the meaning of the formal parameter $M$ in the Schwarzschild solution: namely, we will move to the Newtonian limit. In the context of the equation of motion (18), this implies a lower approximation in the small parameter $2M/b$, within which equation (18) takes the form
$\mathcal{E}=M+\frac{M}{2}\left(\frac{db}{dt}\right)^{2}-\frac{M^{2}}{b}.$ (19)
Let us compare this expression with the result that follows directly from Newton’s theory of the gravitational field for the same spherical wormhole of mass $M$ and radius $b$. This requires calculating the corresponding energy $U$ of the gravitational field, which in the Newtonian approximation is a locally well-determined quantity:
$U=-\frac{1}{8\pi}\int\mathfrak{g}^{2}\,dV.$ (20)
Here, $\mathfrak{g}$ is the local acceleration of gravity, which in the spherically symmetric case has an absolute value
$g=\frac{M}{b^{2}},\quad(r\geq b),$ (21)
and the integration extends to the entire two-sheeted space. This corresponds to a double integration of expression (20) over the region $r\geq b$ of the ordinary one-sheeted space and leads to the following expression for the total gravitational energy:
$U(b)=-\frac{2}{8\pi}\int_{r\geq b}g^{2}\,dV=-\frac{M^{2}}{b}.$ (22)
The factor 2 in front of the integral in this expression appears precisely because of the two-sheeted nature of space and the double integration over the region $r\geq b$ . After this, the Newtonian wormhole Lagrangian, written in the usual form,
$L\!\left(b,\frac{db}{dt}\right)=\frac{M}{2}\left(\frac{db}{dt}\right)^{2}-U(b),$ (23)
immediately leads to the conservation of the total energy of the wormhole, which includes the energy of the field (22):
$\frac{M}{2}\left(\frac{db}{dt}\right)^{2}-\frac{M^{2}}{b}=\text{const.}$ (24)
By denoting the constant on the right-hand side as EME-M, we arrive at the law of conservation of the total energy $E$ of the wormhole:
$E=M+\frac{M}{2}\left(\frac{db}{dt}\right)^{2}-\frac{M^{2}}{b}.$ (25)
This expression for the total energy $E$ coincides with the previous expression (19) for the formal parameter $\mathcal{E}$.
Thus, in the Newtonian limit, the formal conserved quantity $\mathcal{E}$ arising in the equation of motion (18) naturally reproduces the wormhole’s self-energy $E$, which includes the field energy. This provides grounds for maintaining the same interpretation in the relativistic case, where the gravitational field energy is no longer a locally, but rather a globally defined quantity. For this reason, we shall henceforth refer to the corresponding integral of motion $\mathcal{E}$ as the wormhole’s self-energy, without any further qualifications.
We have thus demonstrated that, in the Newtonian limit, the extended formulation of the equivalence principle we adopted is completely consistent with ordinary classical mechanics—both lead to identical equations of motion (19) and (25). It is precisely the correspondence principle that reduces the applicability of this extended formulation to the classical mechanics, which is initially based on our everyday experience. A rigorous derivation of the extended equivalence principle from the polynomial action is a much more difficult task and is left for future research.
Another important observation following from this analysis is the significant influence of the two-sheeted space on the wormhole dynamics. The factor 2 in front of the integral in expression (22) is a direct consequence of the two-sheeted space of the wormhole. In a one-sheeted space, this factor would be equal to 1, which would lead to a half-expression for the field energy $U(b)$. This would cause a discrepancy between the equation of motion for a test particle, determined by the potential M/b-M/b, and the analogous equation of motion for the surface of a matter-free object in a one-sheeted space, determined by the half-potential M/(2b)-M/(2b). In turn, such a discrepancy would result in desynchronization of the motion of the test particle and the gravitating surface (see Fig. 1), and to a violation of the equivalence principle in its extended formulation.
This means that the extended formulation of the equivalence principle not only allows for a two-sheeted structure of space, but also requires it - as a necessary condition for the existence of matter-free sources of the gravitational field.
In this sense, the extended formulation of the equivalence principle and the two-sheetedness of the wormhole space manifest themselves as mutually conditioned factors, which imparts rigidity to their combination in the theory of a degenerate wormhole. Of course, the example of a degenerate wormhole considered here has a limited degree of generality. However, it alone already indicates that, at least in the case of spherical symmetry,
consistency of the equivalence principle with Newtonian mechanics requires a two-sheeted structure.
@@
Note that here the limit $b\rightarrow 2M$ does not produce the usual Schwarzschild interior region $0<r<2M$, but rather the Einstein–Rosen bridge connecting two copies of the exterior Schwarzschild region $r\geq 2M$ via the degenerate throat at $r=2M$. In the Einstein-Rosen-Horowitz-Klinkhamer polynomial formulation, this configuration is geodesically complete and free of curvature singularities on both sheets; the singular region $r=0$ of the standard maximally extended Schwarzschild solution is simply absent from the manifold.
@@
By ‘stable final state’ we mean that the collapsing degenerate wormhole asymptotically approaches this static Einstein–Rosen configuration and cannot evolve further. No further contraction beyond $b=2M$ is possible because the manifold by construction has no region with $r<2M$.
@@
In each of these three regimes, the Klinkhamer wormhole, although non-stationary, nevertheless remains traversable for a certain time. This time depends on the initial conditions and can be quite significant. For example, the collapse from a resting state with an initial radius $b(0)\equiv b_{0}\gg 2M$ takes a time $\tau$ equal to the fall time of a test particle from a distance $b_{0}$ in the gravitational field of a classical point source of mass $M$. This time is (Landau and Lifshitz, 1982), Eq.(103.17):